# Tracing the Techno-Supremacy Doctrine: A Critical Discourse Analysis of the AI Executive Elite

### v1

@@ Abstract
This paper critically analyzes the discourse of the ‘AI executive elite,’ a group of highly influential individuals shaping the way AI is funded, developed, and deployed worldwide. The primary objective is to examine the presence and dynamics of the ‘Techno-Supremacy Doctrine’ (TSD), a term introduced in this study to describe a belief system characterized by an excessive trust in technology’s alleged inherent superiority in solving complex societal problems. This study integrates quantitative heuristics with in-depth qualitative investigations. Its methodology is operationalized in a two-phase critical discourse analysis of 14 texts published by elite members between 2017 and 2025. The findings demonstrate that the elite is not a monolithic bloc but exhibits a broad spectrum of stances. The discourse is highly dynamic, showing a marked polarization and general increase in pro-TSD discourse following the launch of ChatGPT. The analysis identifies key discursive patterns, including a dominant pro-TSD narrative that combines utopian promises with claims of inevitable progress, and the common tactic of acknowledging risks only as a strategic preamble to proposing further technological solutions. This paper presents TSD as a comprehensive analytical framework and provides a ‘diagnostic toolkit’ for identifying its manifestations, from insidious to benign. It argues that fostering critical awareness of these discursive patterns is essential for AI practitioners, policymakers, and the public to actively navigate the future of AI.
Keywords Techno-Optimism, Responsible AI, Critical Discourse Analysis
@@ 1 Introduction
The power of technology is undeniable and its allure is seductive, particularly in light of the seemingly ever-growing capabilities of the latest generative AI models. Amid this awe-inspiring and accelerated development, it is tempting to think of technology as inherently superior to tackle not only mundane tasks, but also complex societal issues. This trust in technology’s prowess can become excessive and uncritical, constituting the foundation of the Techno-Supremacy Doctrine (TSD).
@@
Adhering to TSD can lead to oversimplifying problems, devaluating alternative approaches, and uncritically accepting technology. For example, TSD can manifest perniciously in the field of education, where the intricate process of learning is often reduced to a quantifiable output amenable to optimization. This mindset posits that systemic issues like educational disparity can be “fixed” by deploying technological tools (e.g., laptops, AI tutors). While the use of AI as an aid for education is promising, there are significant risks, as it may affect cognitive processes that are essential for genuine learning. Recent research shows that students using Large Language Models (LLMs) for essay writing displayed significantly weaker brain engagement, accumulated a “cognitive debt,” and felt a lower sense of ownership over their work compared to control groups (Kosmyna et al., 2025). Focusing on the final product (e.g., an essay) over the cognitive effort can narrow educational goals to “teaching to the test,” substituting deep understanding and expert human mentorship for the optimization of simplistic metrics (Toyama, 2015, p. 94).
@@
TSD’s risks are significantly amplified when it is adopted by people with the power to direct how AI is funded, developed, and deployed on a global scale. Through their discourse, this AI executive elite uses public platforms to shape narratives, influence policy, and define what is considered possible, desirable, and inevitable (e.g., AI tutors). The dramatic acceleration in AI development following the launch of ChatGPT has been accompanied by a recent shift in global discourse, from an initial focus on safety and ethics to one that champions rapid action and innovation. This AI ‘gold rush’ makes it more urgent than ever to develop a critical sensitivity for TSD. For AI to be developed responsibly and democratically, stakeholders must transition from a passive reception of TSD’s alluring narratives to a state of active, critical engagement. This is the core rationale for this paper, which is based on the premise that for TSD to be effectively challenged, it must first be systematically identified and understood.
@@
The primary aim of this paper is to critically analyze the discourse of the AI executive elite to understand the presence, function, and dynamics of the Techno-Supremacy Doctrine. This research is guided by three interconnected objectives:
  • To critically analyze whether and how aspects of TSD are presented, challenged, or reproduced (O1).
  • To critically analyze the discursive strategies used to respond to AI-related concerns (O2).
  • To critically analyze how this discourse has changed over time and provide potential explanations in terms of its evolving broader socio-political context (O3).
This study’s contributions can be grouped into three main categories: theoretical, methodological, and empirical.
@@
Theoretically, while other studies have examined individual ideologies, this paper articulates the Techno-Supremacy Doctrine as a comprehensive, synthesized framework. By integrating four previously distinct belief systems (i.e., Tech Goggles, Solutionism, Techno-chauvinism, and the Cult of Technology) into a single, self-reinforcing cyclical model, the TSD framework provides a more holistic analytical lens than studying these mindsets in isolation. It offers a structured way to understand how technical framing, solutionist impulses, chauvinistic beliefs, and quasi-religious faith can interlink to form a resilient and robust worldview. Furthermore, the study introduces the conceptual distinction of Benign Techno-Optimism (BTO). This concept provides a crucial nuance to the study of techno-optimist discourse, moving beyond simplistic dichotomies (e.g., pro- versus anti-TSD, techno-optimism versus techno-pessimism). It allows for a more fine-grained analysis that can differentiate between uncritical, pernicious forms of TSD and more self-aware, critically tempered expressions of optimism, providing a richer understanding of the spectrum of beliefs within the elite.
@@
Methodologically, in a field often characterized by high-level commentary, this study develops and demonstrates a practical and scalable two-phase approach to Critical Discourse Analysis that systematically balances the need for a broad, corpus-wide overview with the necessity of in-depth qualitative analysis of specific phenomena. Additionally, this paper introduces metrics as novel quantitative heuristics. By translating detailed qualitative codings into transparent, comparable scores, these metrics provide a systematic way to identify patterns, variations, and outliers, thereby grounding the selection of texts for deeper analysis in verifiable data. This adaptation enhances the rigor and transparency of the analytical process, responding directly to scholarly critiques of CDA concerning potential analyst bias and a lack of systematic procedure.
@@
Empirically, rather than focusing on cherry-picked pronouncements typical in media reporting, this research provides a detailed CDA of the public discourse of a systematically defined elite during the pivotal 2017-2025 period. It offers concrete, linguistically-grounded evidence of the spectrum of TSD adherence within this highly influential group, challenging monolithic portrayals of “Silicon Valley thinking.” The findings reveal the dynamic, adaptive nature of this elite discourse, including a marked polarization of stances in the post-ChatGPT era and significant shifts within the narratives of individual leaders. By tracing these dynamics and linking them to a detailed timeline of technological, political, and social events, this study provides a crucial, evidence-based account of how the struggle to define the future of AI is being waged discursively by some of its most powerful proponents.
@@ 2 The Techno-Supremacy Doctrine
TSD is a belief system characterized by an excessive and often uncritical trust in technology’s alleged inherent superiority to address complex societal challenges. TSD represents a mindset that overvalues technology’s capacity to solve problems and often underestimates the complexity of social issues, as well as the importance of human agency, values, and context. It fosters the idea that technological advancements alone are sufficient for societal progress, potentially leading to the neglect of non-technological solutions and a blindness to the limitations and potential negative consequences of purely technological approaches.
The term ‘Techno-Supremacy Doctrine’ was selected over several alternatives (e.g., Techno-Primacy Principle, Algorithmic Solutionist Worldview, Techno-Centric Goggles) for its precision in capturing this belief system’s distinct ideological nature. The prefix ‘Techno’ grounds the concept firmly in the domain of technology in its broadest sense, encompassing not just AI but technology in general. The core component, ‘Supremacy,’ denotes more than mere primacy. It highlights a conviction in technology’s inherent dominance, ultimate authority, and fundamental superiority over alternative non-technological approaches to problem-solving. This term reflects the overvaluation and uncritical trust central to TSD, where technological considerations often override or marginalize other vital human and contextual factors. Finally, ‘Doctrine’ is employed because TSD represents not an isolated bias but a systemic and self-reinforcing set of interconnected beliefs, assumptions, and guiding tenets that shape perception and action. It is important to note that ‘Doctrine’ is used as an analytical construct to describe the function and logic of this belief system, rather than a claim that its tenets are always consciously held by its adherents.
@@
  1. 1.
    Tech Goggles (Green, 2019). Framing a wide array of societal issues primarily as technology problems (e.g., fairness being reduced to a quantitative metric). Given a problem, Tech Goggles adherents are inclined to focus on determining how technology can address it.
  2. 2.
    Solutionism (Morozov, 2013). Regarding complex social, political, and human issues as problems with singular, often technological, solutions or "fixes" (e.g., believing education disparity can be primarily addressed with AI tutors). Solutionism often leads to a drive to implement solutions rapidly and widely.
  3. 3.
    Techno-Chauvinism (Broussard, 2018). Considering technology as inherently superior or objective. Techno-Chauvinism adherents are firmly convinced that technology is the optimal solution to any problem, including complex societal issues, regardless of the context (i.e., believing there is no problem technology cannot fix).
  4. 4.
    Cult of Technology (Toyama, 2015). Having a quasi-religious confidence in technology’s alleged inherent power. Cult of Technology adherents are often overly optimistic about the future based on assumed, almost magical, capabilities of current or future technologies.
@@
TSD’s components contribute to a cycle of mutual reinforcement and amplification (cf. Figure 1). Tech Goggles serves as an entry point, predisposing one to perceive and frame complex societal issues primarily through a technological lens. This framing naturally leads to Solutionism, narrowing the focus onto identifying singular, often technological “solutions” or “fixes.” Techno-Chauvinism provides justification, asserting the inherent superiority of the chosen approach over non-technological alternatives. The perceived effectiveness of this solution then fuels the Cult of Technology, strengthening the underlying faith in technology’s power. This reinforced faith, in turn, makes the individual more inclined to habitually adopt Tech Goggles when facing future challenges, thereby perpetuating the cycle.
@@ 2.1 Related Concepts
At its foundation, TSD is based on the premise of Technological Determinism (Dotson, 2015; Héder, 2021), the belief that technology, often portrayed as an inevitable and unstoppable force, is the primary driver of social change (e.g., attributing the spread of literacy primarily to the invention of the printing press). TSD expands on this deterministic view, adding a normative dimension: technology-driven change is inherently progressive, and technology is superior to all other problem-solving strategies. It elevates the methods of computational thinking (Wing, 2006) into a worldview, positing this approach as the optimal way to address complex societal problems. While a form of techno-optimism, TSD goes beyond a mere positive outlook to a doctrine of inherent superiority, sometimes manifesting as the quasi-religious faith seen in the Cult of Technology component.
@@
It is common for techno-optimists to assert that technology is neither good nor bad, but a value-neutral tool (cf. Value-Neutrality Thesis (VNT) (Miller, 2021)). TSD neither upholds nor denies VNT, as it merely claims technology’s inherent instrumental (as opposed to moral) superiority. However, by framing technology as uniquely powerful and scalable, but ultimately value-neutral, TSD proponents can deflect responsibility (cf. “misuse,” “bad actors”) and portray risks as merely problems to be managed by better (e.g., more secure) technology, thereby reinforcing core TSD components like Tech Goggles and Solutionism.
@@ 2.2 Concerns
Problem oversimplification arises significantly from the interaction of Tech Goggles and Solutionism. As critiqued by Green and Morozov, this combination leads to framing diverse societal challenges primarily as narrow problems amenable to singular, often technological, solutions. This framing can lead to a rush to adopt such solutions rapidly and widely before fully understanding the issue’s nuances or even properly defining them (i.e., reaching for the answer before the question has been fully asked). This perspective inherently conflicts with deep contextual understanding and holistic problem assessment before intervention. It often oversimplifies or ignores the historical, social, economic, and political dimensions of societal issues that non-technological ethical frameworks prioritize.
Techno-Chauvinism strongly drives the devaluation of alternative approaches. As Broussard critiques, this mindset inherently considers technology superior and more objective, thereby devaluing non-technological approaches as inferior (e.g., sub-optimal, inefficient, subjective). This stance directly opposes calls for critical evaluation of whether technology is the most appropriate approach, leading to the neglect of alternative methods that might be more effective or ethically sound within a given social and political context. Moreover, this chauvinism often devalues situated knowledge from users and other affected stakeholders during technology design (unlike approaches like Participatory Design (Schuler and Namioka, 1993)).
The Cult of Technology fuels the uncritical acceptance of technology. Toyama’s critique highlights the manifestation of this acceptance as an almost religious conviction in technology’s alleged inherent transformative power, based on an overestimation of current or future capabilities. This blind faith discourages skepticism and rigorous testing, which is central to responsible innovation, often resulting in the marginalization of potential negative consequences and the adoption of solutions without sufficient scrutiny, simply because they are technologically advanced. This stance contrasts with acknowledging that technology’s impact is contingent on human agency and socio-institutional contexts, rather than an intrinsic magical quality.
An overarching concern about TSD as a whole is that its adoption by people holding social, political, or economic power (e.g., the AI executive elite) directly undermines efforts towards responsible innovation. Adherence to TSD by influential people can lead to prioritizing technology research over other critical areas. Investments in social programs, policy reforms, ethical oversight, and human capital can be neglected, thereby impeding a holistic and ethical approach to societal progress.
@@
3 Methodology
3.1 Data Analysis

### v2

@@ Abstract
This paper critically analyzes the discourse of the ‘AI executive elite,’ a group of highly influential individuals shaping the way AI is funded, developed, and deployed worldwide. Its primary objective is to examine the presence and dynamics of the ‘Techno-Supremacy Doctrine’ (TSD), a term coined for this research to describe a belief system characterized by an excessive trust in technology’s alleged inherent superiority in solving complex societal problems. This paper presents a novel approach to Critical Discourse Analysis (CDA) that responds to scholarly critiques by integrating quantitative metrics with in-depth qualitative investigations. This methodology is operationalized to analyze 14 texts published by elite members between 2017 and 2025. The findings demonstrate that the elite is not a monolithic bloc but exhibits a broad spectrum of stances. The discourse is highly dynamic, showing a marked polarization and general increase in pro-TSD discourse following the launch of ChatGPT. The analysis identifies key discursive patterns, including a dominant pro-TSD narrative that combines utopian promises with claims of inevitable progress, and the common tactic of acknowledging risks only as a strategic preamble to proposing further technological solutions. This approach to CDA acts as a linguistic ‘diagnostic toolkit’ for identifying TSD’s diverse manifestations, from insidious to benign. This paper argues that fostering critical awareness of TSD’s discursive patterns is essential for AI practitioners, policymakers, and the general public to actively navigate the future of AI.
Keywords Techno-Supremacy Doctrine, Techno-Optimism, Artificial Intelligence, Critical Discourse Analysis
@@ 1 Introduction
The power of technology is undeniable, and its allure seductive, particularly in light of the seemingly ever-growing capabilities of the latest generative Artificial Intelligence models (e.g., ChatGPT). Amid this admittedly awe-inspiring and accelerated development, it is tempting to think of technology as inherently superior at tackling not only mundane tasks but also complex societal issues. This trust in technology’s prowess can become excessive and uncritical, constituting the foundation of the Techno-Supremacy Doctrine (TSD).
@@
Adhering to TSD can be problematic. It can lead to oversimplifying problems, devaluing alternative approaches, and uncritically accepting technology. For example, TSD can manifest perniciously in education, where the intricate process of learning is often reduced to a quantifiable output amenable to optimization. This mindset posits that systemic issues such as educational disparities can be “fixed” by deploying technological tools (e.g., laptops, AI tutors). While the use of AI as an educational aid is promising, there are significant risks, as it may affect cognitive processes essential to genuine learning. Recent research shows that students using Large Language Models (LLMs) for essay writing displayed significantly weaker brain engagement, accumulated a “cognitive debt,” and felt a lower sense of ownership over their work compared to control groups (Kosmyna et al., 2025). Focusing on the final product (e.g., an essay) over the cognitive effort can narrow educational goals to “teaching to the test,” substituting deep understanding and expert human mentorship for the optimization of simplistic metrics (Toyama, 2015, p. 94).
@@
TSD’s risks are significantly amplified when it is adopted by people with the power to direct how AI is funded, developed, and deployed on a global scale. Through their discourse, this AI executive elite uses public platforms to shape narratives, influence policy, and define what is considered possible, desirable, and inevitable (e.g., AI tutors). The dramatic acceleration in AI development following the launch of ChatGPT has been accompanied by a recent shift in global discourse, from an initial focus on safety and ethics to one that champions rapid action and innovation. This AI ‘gold rush’ makes it more urgent than ever to develop a critical sensitivity to TSD. For AI to be developed responsibly and democratically, stakeholders must transition from passive reception of TSD’s alluring narratives to active, critical engagement. This is the core rationale for this paper, which is based on the premise that for TSD to be effectively challenged, it must be systematically understood and identified.
@@
The primary aim of this paper is to critically analyze the discourse of the AI executive elite to understand the presence, function, and dynamics of the Techno-Supremacy Doctrine. This research is guided by three objectives: (I) to analyze whether and how aspects of TSD are presented, challenged, or reproduced; (II) to identify which discursive strategies are used to respond to AI-related concerns; and (III) to investigate how the elite’s discourse has changed over time and provide potential explanations in terms of its evolving broader socio-political context.
This paper offers theoretical, methodological, and empirical contributions.
@@
Theoretically, this paper articulates TSD as a comprehensive, synthesized theoretical framework that builds on established belief systems. It offers a structured way to understand how technical framing, solutionist impulses, chauvinistic beliefs, and quasi-religious faith can interlink to form a resilient, robust, and self-reinforcing worldview. Moreover, the concept of Benign Techno-Optimism is introduced as a particular incarnation of TSD. This concept provides a crucial nuance to the study of techno-optimist discourse, moving beyond simplistic dichotomies (e.g., techno-optimism versus techno-pessimism). It allows for a more fine-grained analysis that can differentiate between uncritical, pernicious forms of TSD and more self-aware, critically tempered expressions of techno-optimism, providing a richer understanding of the spectrum of beliefs within the elite.
@@
Methodologically, in a field often characterized by high-level commentary, this paper develops a two-phase ‘T’ approach to Critical Discourse Analysis (CDA) that systematically balances the need for a broad, corpus-wide overview with the necessity of in-depth qualitative analysis of specific phenomena. Novel metrics are introduced as quantitative heuristics. By translating detailed qualitative codings into transparent, comparable scores, these metrics provide a systematic way to identify patterns, variations, and outliers, thereby grounding the selection of texts for deeper analysis in verifiable data. This adaptation enhances the rigor and transparency of the analytical process, directly responding to scholarly critiques of CDA regarding potential analyst bias and a lack of systematic procedure.
@@
Empirically, this paper employs CDA approach mentioned previously to analyze the public discourse of a systematically defined elite between 2017 and 2025. The analysis offers concrete, linguistically-grounded evidence of the spectrum of TSD adherence within this highly influential group, challenging monolithic portrayals of “Silicon Valley thinking.” The findings reveal the dynamic, adaptive nature of this elite discourse, including a marked polarization of stances in the post-ChatGPT era and significant shifts within the narratives of individual leaders. By tracing these dynamics and linking them to a detailed timeline of technological, political, and social events, this paper provides a crucial, evidence-based account of how the struggle to define the future of AI is being waged discursively by some of its most powerful proponents.
After presenting TSD as the study’s theoretical foundation, the paper details the ‘T’ approach to CDA developed for this research. It then describes the empirical study by first outlining the criteria for identifying the AI executive elite and selecting the research corpus, and then presenting findings in terms of TSD’s discursive spectrum and temporal dynamics. Key results and important considerations are then discussed, followed by a conclusion that highlights limitations and avenues for future research.
@@ 2 The Techno-Supremacy Doctrine
The Techno-Supremacy Doctrine is a belief system characterized by excessive, often uncritical trust in technology’s alleged inherent superiority to address complex societal challenges. TSD represents a mindset that overvalues technology’s capacity to solve problems and often underestimates the complexity of social issues and the importance of human agency, values, and context. It fosters the idea that technological advancements alone are sufficient for societal progress, potentially leading to the neglect of non-technological solutions and a blindness to the limitations and potential negative consequences of purely technological approaches.
The term ‘Techno-Supremacy Doctrine’ was selected over several alternatives for its precision in capturing this belief system’s distinct ideological nature. The prefix ‘Techno’ grounds the concept firmly in the domain of technology in its broadest sense, encompassing not just AI but technology in general. The core component, ‘Supremacy,’ denotes more than mere primacy. It highlights a conviction in technology’s inherent dominance, ultimate authority, and fundamental superiority over alternative non-technological approaches to problem-solving. This term reflects the overvaluation and uncritical trust central to TSD, where technological considerations often override or marginalize other vital human and contextual factors. Finally, ‘Doctrine’ is employed because TSD represents not an isolated bias but a systemic and self-reinforcing set of interconnected beliefs, assumptions, and guiding tenets that shape perception and action. Note, however, that we do not claim that its tenets are always consciously held by its adherents.
@@
  • Tech Goggles (Green, 2019): Framing a wide array of societal issues primarily as technology problems (e.g., reducing fairness to a quantitative metric). Given a problem, Tech Goggles adherents tend to focus on how technology can address it.
  • Solutionism (Morozov, 2013): Regarding complex social, political, and human issues as problems with singular, often technological, solutions or “fixes” (e.g., believing education disparity can be primarily addressed with AI tutors). Solutionism often drives the rapid, widespread implementation of ‘solutions.’
  • Techno-Chauvinism (Broussard, 2018): Considering technology as inherently superior or objective. Techno-Chauvinists are firmly convinced that technology is the optimal solution to any problem, including complex societal issues, regardless of context (i.e., believing there is no problem technology cannot fix).
  • Cult of Technology (Toyama, 2015): Having a quasi-religious confidence in technology’s alleged inherent power. Cult of Technology adherents are often overly optimistic about the future based on assumed, almost magical, capabilities of current or future technologies.
@@
TSD’s components contribute to a cycle of mutual reinforcement and amplification (cf. Figure 1). Tech Goggles serves as an entry point, predisposing one to perceive and frame complex societal issues primarily through a technological lens. This framing naturally leads to Solutionism, narrowing the focus onto identifying singular technological “solutions” or “fixes.” Techno-Chauvinism provides justification, asserting the inherent superiority of the chosen approach over non-technological alternatives. The perceived effectiveness of this solution then fuels the Cult of Technology, strengthening the underlying faith in technology’s power. This reinforced faith, in turn, makes the individual more inclined to habitually adopt Tech Goggles when facing future challenges, thereby perpetuating the cycle.
@@ 2.1 Related Concepts
At its foundation, TSD is based on the premise of Technological Determinism (Dotson, 2015; Héder, 2021), the belief that technology, often portrayed as an inevitable and unstoppable force, is the primary driver of social change (e.g., attributing the spread of literacy primarily to the invention of the printing press). TSD expands on this deterministic view, adding a normative dimension: technology-driven change is inherently progressive and superior to all other problem-solving strategies. It elevates the methods of computational thinking (Wing, 2006) into a worldview, positing this approach as the optimal way to address complex societal problems. While a form of techno-optimism, TSD goes beyond a mere positive outlook to a doctrine of inherent superiority, sometimes manifesting as the quasi-religious faith seen in the Cult of Technology component.
@@
It is common for techno-optimists to assert that technology is neither good nor bad, but a value-neutral tool (cf. Value-Neutrality Thesis (VNT) (Miller, 2021)). TSD neither upholds nor denies VNT, as it merely claims technology’s inherent instrumental (rather than moral) superiority. However, by framing technology as uniquely powerful and scalable, but ultimately value-neutral, TSD proponents can deflect responsibility (cf. “misuse,” “bad actors”) and portray risks as merely problems to be managed by better (e.g., more secure) technology, thereby reinforcing core TSD components like Tech Goggles and Solutionism.
@@ 2.2 Concerns
Problem oversimplification arises significantly from the interaction of Tech Goggles and Solutionism. As Green and Morozov critique, this combination leads to framing diverse societal challenges primarily as narrow problems amenable to singular technological solutions. This framing can lead to a rush to adopt such solutions rapidly and widely before fully understanding the issue’s nuances or even properly defining them (i.e., reaching for the answer before the question has been fully asked). This perspective inherently conflicts with deep contextual understanding and holistic problem assessment before intervention. It often oversimplifies or ignores the historical, social, economic, and political dimensions of societal issues that non-technological ethical frameworks prioritize.
Techno-Chauvinism strongly drives the devaluation of alternative approaches. As Broussard critiques, this mindset inherently considers technology superior and more objective, thereby devaluing non-technological approaches as inferior (e.g., sub-optimal, inefficient, subjective). This stance directly opposes calls for critical evaluation of whether technology is the most appropriate approach, thereby neglecting alternative methods that might be more effective or ethically sound within a given social and political context. Moreover, this chauvinism often devalues situated knowledges (Haraway, 1988) from users and other affected stakeholders during technology design (unlike approaches like Participatory Design (Schuler and Namioka, 1993)).
The Cult of Technology fuels the uncritical acceptance of technology. Toyama’s critique highlights the manifestation of this acceptance as an almost religious conviction in technology’s alleged inherent transformative power, based on an overestimation of current or future capabilities. This blind faith discourages skepticism and rigorous testing, which are central to responsible innovation, often leading to the marginalization of potential negative consequences and the adoption of solutions without sufficient scrutiny, simply because they are technologically advanced. This stance contrasts with acknowledging that technology’s impact is contingent on human agency and socio-institutional contexts, rather than an intrinsic magical quality.
An overarching concern about TSD as a whole is that its adoption by people holding social, political, or economic power (e.g., the AI executive elite) directly undermines efforts towards responsible innovation. Adherence to TSD by influential people can lead to prioritizing technology research over other critical areas. Investments in social programs, policy reforms, ethical oversight, and human capital can be neglected, thereby hindering a holistic, ethical approach to societal progress.
@@
3 The ‘T’ Approach to Critical Discourse Analysis